# More is Less: The Pitfalls of Multi-Model Synthetic Preference Data in DPO Safety Alignment

### v1

@@ B.1 SFT Training Hyper Parameters
For our supervised fine-tuning (SFT) stage, we utilized 50k samples from HuggingFaceH4/ultrafeedback_binarized [9] for general capabilities and 50k samples from PKU-Alignment/PKU-SafeRLHF [10] for safety. All experiments were conducted on 4 A100 GPUs. We trained all models for one epoch with the detailed hyperparameters listed in Table 5.
@@ B.2 DPO Training Hyperparameters
B.3 Composition of Multi-Model Response Pool
@@
Figure 7: Composition of chosen and rejected responses from all models pool. “Stronger" here means larger size model (Llama2-70B-Chat, Llama3.1-70B-Instruct, Mixtral-8x7B-Instruct-v0.1 , Qwen2.5-70B-Instruct) compared to the target 7B models
B.4 Safety Evaluation Instructions
We use GPT-4o to perform the evaluation, on the 100 first examples from AdvBench. For the scoring guidelines, we follow [34], which is a revised version from [35]. In our prompt, we include Meta’s usage guidelines. [^1] https://ai.meta.com/llama/use-policy/

### v2

@@ B.1 SFT Training Hyper Parameters
For our supervised fine-tuning (SFT) stage, we utilized 50k samples from HuggingFaceH4/ultrafeedback_binarized (Cui et al., 2024) for general capabilities and 50k samples from PKU-Alignment/PKU-SafeRLHF (Dai et al., 2023) for safety. All experiments were conducted on 4 A100 GPUs. We trained all models for one epoch with the detailed hyperparameters listed in Table 9.
@@ B.2 DPO Training Hyperparameters
B.3 DPO Parameter Ablation
To isolate the cause of reward hacking discussed in Section  3.4, we conducted controlled experiments to determine whether this phenomenon results solely from the data or could be mitigated through hyperparameter adjustments. We ran experiments on the GPT-4o + Self dataset using Llama3.1-8B, systematically varying $\beta$ from 0.01 to 0.9 and testing two learning rates: 5e-7 and 1e-7.
Table 11 presents the Attack Success Rate (ASR) results across all parameter combinations. As shown, ASR remained consistently high across all settings, ranging from 0.53 to 0.57. This consistent performance across different hyperparameter configurations suggests that hyperparameter tuning does not alleviate reward hacking, reinforcing our conclusion that the issue stems from the data itself.
Table 11: DPO Parameter Ablation: Attack Success Rate across different $\beta$ and learning rate combinations
$\beta=0.01$ $\beta=0.1$ $\beta=0.2$ $\beta=0.3$ $\beta=0.4$ $\beta=0.5$ $\beta=0.6$ $\beta=0.7$ $\beta=0.8$ $\beta=0.9$
$lr=1e-7$ 0.54 0.54 0.56 0.56 0.53 0.53 0.56 0.56 0.56 0.57
$lr=5e-7$ 0.55 0.56 0.55 0.56 0.55 0.56 0.54 0.53 0.54 0.56
B.4 Composition of Multi-Model Response Pool
@@
Figure 7: Composition of chosen and rejected responses from all models pool. “Stronger” here means larger size model (Llama2-70B-Chat, Llama3.1-70B-Instruct, Mixtral-8x7B-Instruct-v0.1 , Qwen2.5-70B-Instruct) compared to the target 7B models
B.5 Safety Evaluation Instructions
We use GPT-4o to perform the evaluation, on the 100 first examples from AdvBench. For the scoring guidelines, we follow (He et al., 2024), which is a revised version from (Qi et al., 2023). In our prompt, we include Meta’s usage guidelines. [^2] https://ai.meta.com/llama/use-policy/